OpenStudy (anonymous):

a country's population in 1990 was 134 million in 2000 it was 139 million estimate the population in 2014

OpenStudy (anonymous):

HI :DDD

OpenStudy (anonymous):

Subtract 139 million and 134 million. Then divide by 134 million. This will give you a percent that you will multiply by 139 million...this will give the growth...add it to 139 million

OpenStudy (anonymous):

i got 144,186,567 so 144million?

OpenStudy (anonymous):

139 - 134 million = 5,000,000/139,000,000=3.6% 139 million * 0.036 =5004000 139 million + 5,004,000=144,004,000

OpenStudy (anonymous):

yes